what are the Newton equations of motion for constant acceleration.
1
Expert's answer
2020-09-14T10:22:29-0400

We know that in an inertial reference frame the motion follows an equation

"ma = m\cdot \dfrac{d^2r}{dt^2} = F."

The constant acceleration means "\dfrac{d^2r}{dt^2} =a" , so

"v(t) = \dfrac{dr}{dt} = at+v_0"

and "r(t) = a\dfrac{t^2}{2} + v_0t + r_0."

Here "v_0" is the initial velocity and "r_0" is the initial position.

The last equation can be rewritten in terms of the distance: "s(t) = a\dfrac{t^2}{2} + v_0t" .

Here (https://thecuriousastronomer.wordpress.com/2014/11/11/newtons-equations-of-motion-revisited/) the author presents three additional equations ([3], [4], [5]), that can be obtained by rewriting the equation above in terms of the velocity after time t (v(t)). The most important of the three is [4], because it helps us to link the velocity and the distance without the time.